# Chem-FINESE: Validating Fine-Grained Few-shot Entity Extraction through Text Reconstruction

### v3

@@ 3.2 Entity Extraction Module
$\mathcal{L}_{\mathrm{gen}}=\sum_{t=1}^{T}\log p(y_{t}|\mathcal{S},y_{<t}).$ (2)
@@ 3.3 Self-validation Module
$\mathcal{L}_{\mathrm{recon}}=\sum_{\hat{t}=1}^{\hat{T}}\log p(\hat{s}_{\hat{t}}|\mathbf{H},\hat{s}_{<\hat{t}}),$ (4)
@@ 3.5 Training Objective
Dataset Split #Pair
#$\mathbf{\overline{\#Token}}$
#$\mathbf{\overline{\#Entity}}$
Train
542
32.9
3.10
ChemNER+
Valid
100
39.9
4.57
Test
100
39.4
4.61
@@
Train
6,561
37.8
1.57
CHEMET
Valid
520
31.6
2.15
Test
663
36.6
1.95
@@ 4.1 Dataset Creation
$k$-shot
6 9 12 15 18
RoBERTa
8.09
7.98
8.00
16.22
7.94
PubMedBERT
5.48
5.12
5.77
5.46
5.88
ScholarBERT
23.96
29.82
27.65
31.48
32.76
@@
$k$-shot
6 9 12 15 18
RoBERTa
4.91
4.16
4.79
4.83
4.81
PubMedBERT
4.07
4.67
3.87
4.47
3.96
ScholarBERT
17.00
33.63
29.65
29.72
32.52
@@ 5.2 Overall Performance
$k$-shot
6 9 12 15 18
InBoXBART
36.96
38.22
38.34
47.91
42.84
+ Valid
45.07
45.28
41.56
48.15
46.15
+ Valid + CL
45.58
44.03
45.25 51.68 47.88
@@
$k$-shot
6 9 12 15 18
InBoXBART
46.74
42.07
44.32
47.58
52.90
+ Valid
47.87
46.01
44.18
50.55
50.50
+ Valid + CL
48.96 49.83 47.03 50.61 54.10
@@
$k$-shot
6 9 12 15 18
RoBERTa
2.04
2.05
2.05
0.00
2.05
PubMedBERT
2.05
0.00
0.00
2.13
0.00
ScholarBERT
0.00
9.28
4.71
0.00
6.90
@@
$k$-shot
6 9 12 15 18
RoBERTa
0.00
0.00
0.00
0.00
0.00
PubMedBERT
0.00
0.00
0.00
0.00
0.00
ScholarBERT
0.00
0.00
0.00
0.00
0.00
@@ 6.1 Qualitative Analysis
InBoXBART
Several cyclophanes, polycycles, … have been synthesized by employing a combination of Suzuki cross-coupling and metathesis Coupling reactions.
@@
+Valid
Several cyclophanes Heterocyclic compounds, polycycles, … have been synthesized by employing a combination of Suzuki cross-coupling and metathesis Organic reactions.
@@
+Valid+CL
Several cyclophanes Heterocyclic compounds, polycycles Biomolecules, … have been synthesized by employing a combination of Suzuki cross-coupling Coupling reactions and metathesis Chemical properties.
@@
Ground Truth
Several cyclophanes Aromatic compounds, polycycles Organic polymers, … have been synthesized by employing a combination of Suzuki cross-coupling Coupling reactions and metathesis Substitution reactions.
InBoXBART
… with the advantages of asymmetric catalysis (step and atom economy) in a rare example of an enantioselective cross coupling of a racemic electrophile bearing an oxygen leaving group Catalysis … the identification of a highly enantioselective process.
@@
+Valid
… with the advantages of asymmetric catalysis (step and atom economy) in a rare example of an enantioselective cross coupling of a racemic electrophile bearing an oxygen leaving group Organometallic compounds … the identification of a highly enantioselective process
@@
+Valid+CL
…with the advantages of asymmetric catalysis Catalysis (step and atom economy) in a rare example of an enantioselective cross coupling of a racemic electrophile bearing an oxygen leaving group Functional groups … the identification of a highly enantioselective process Chemical properties.
@@
Ground Truth
… with the advantages of asymmetric catalysis Catalysis ( step and atom economy ) in a rare example of an enantioselective cross coupling Coupling reactions of a racemic electrophile Organic compounds bearing an oxygen leaving group Functional groups … the identification of a highly enantioselective process Catalysis.
@@ 6.2 Compatible with Other Few-shot Datasets?
Dom. Train Valid Test #Type
#$\mathbf{\overline{\#Token}}$
#$\mathbf{\overline{\#Entity}}$
AI
100
350
430
14
31.5
4.42
Lit.
99
400
416
12
37.6
5.39
Mus.
100
380
465
13
41.4
7.05
Pol.
200
541
651
9
43.5
6.46
Sci.
200
450
543
17
35.8
5.62
@@
BART-base
59.20
66.90
62.78
67.99
62.18
+ Valid
61.84
67.97
60.94
67.22
62.40
+ Valid + CL
62.48 68.22 63.39 68.03 62.87
@@ Appendix A Training and Evaluation Details
RoBERTa
16min
125M
PubMedBERT
18min
109M
ScholarBERT
19min
355M
InBoXBART
58min
139M
+Valid
56min
279M
+Valid+CL
59min
279M
@@ Appendix C Evaluation on Whole Dataset
In-BoXBART
55.73
43.28
48.72
+ Valid
57.49
45.77
50.97
+ Valid + CL
57.41
46.20 51.10
@@
In-BoXBART
64.94
41.62
50.73
+ Valid
70.09 42.16 52.65
+ Valid + CL
68.50
41.31
51.15

### v4

@@ 3.2 Entity Extraction Module
$\mathcal{L}_{\mathrm{gen}}=-\sum_{t=1}^{T}\log p(y_{t}|\mathcal{S},y_{<t}).$ (2)
@@ 3.3 Self-validation Module
$\mathcal{L}_{\mathrm{recon}}=-\sum_{\hat{t}=1}^{\hat{T}}\log p(\hat{s}_{\hat{t}}|\mathbf{H},\hat{s}_{<\hat{t}}),$ (4)
@@ 3.5 Training Objective
Dataset Split #Pair #$\mathbf{\overline{\#Token}}$$\mathbf{\overline{\#Entity}}$
Train 542 32.9 3.10
ChemNER+ Valid 100 39.9 4.57
Test 100 39.4 4.61
@@
CHEMET Valid 520 31.6 2.15
Test 663 36.6 1.95
@@ 4.1 Dataset Creation
$k$-shot 6 9 12 15 18
RoBERTa 8.09 7.98 8.00 16.22 7.94
PubMedBERT 5.48 5.12 5.77 5.46 5.88
ScholarBERT 23.96 29.82 27.65 31.48 32.76
@@
$k$-shot 6 9 12 15 18
RoBERTa 4.91 4.16 4.79 4.83 4.81
PubMedBERT 4.07 4.67 3.87 4.47 3.96
ScholarBERT 17.00 33.63 29.65 29.72 32.52
@@ 5.2 Overall Performance
$k$-shot 6 9 12 15 18
InBoXBART 36.96 38.22 38.34 47.91 42.84
+ Valid 45.07 45.28 41.56 48.15 46.15
+ Valid + CL 45.58 44.03 45.25 51.68 47.88
@@
$k$-shot 6 9 12 15 18
InBoXBART 46.74 42.07 44.32 47.58 52.90
+ Valid 47.87 46.01 44.18 50.55 50.50
+ Valid + CL 48.96 49.83 47.03 50.61 54.10
@@
$k$-shot 6 9 12 15 18
RoBERTa 2.04 2.05 2.05 0.00 2.05
PubMedBERT 2.05 0.00 0.00 2.13 0.00
ScholarBERT 0.00 9.28 4.71 0.00 6.90
@@
$k$-shot 6 9 12 15 18
RoBERTa 0.00 0.00 0.00 0.00 0.00
PubMedBERT 0.00 0.00 0.00 0.00 0.00
ScholarBERT 0.00 0.00 0.00 0.00 0.00
@@ 6.1 Qualitative Analysis
InBoXBART Several cyclophanes, polycycles, … have been synthesized by employing a combination of Suzuki cross-coupling and metathesis Coupling reactions.
@@
InBoXBART … with the advantages of asymmetric catalysis (step and atom economy) in a rare example of an enantioselective cross coupling of a racemic electrophile bearing an oxygen leaving group Catalysis … the identification of a highly enantioselective process.
@@ 6.2 Compatible with Other Few-shot Datasets?
Dom. Train Valid Test #Type #$\mathbf{\overline{\#Token}}$$\mathbf{\overline{\#Entity}}$
AI 100 350 430 14 31.5 4.42
Lit. 99 400 416 12 37.6 5.39
Mus. 100 380 465 13 41.4 7.05
Pol. 200 541 651 9 43.5 6.46
Sci. 200 450 543 17 35.8 5.62
@@
+ Valid 61.84 67.97 60.94 67.22 62.40
+ Valid + CL 62.48 68.22 63.39 68.03 62.87
@@ Appendix A Training and Evaluation Details
RoBERTa 16min 125M
PubMedBERT 18min 109M
ScholarBERT 19min 355M
InBoXBART 58min 139M
+Valid 56min 279M
+Valid+CL 59min 279M
@@ Appendix C Evaluation on Whole Dataset
In-BoXBART 55.73 43.28 48.72
+ Valid 57.49 45.77 50.97
+ Valid + CL 57.41 46.20 51.10
@@
In-BoXBART 64.94 41.62 50.73
+ Valid 70.09 42.16 52.65
+ Valid + CL 68.50 41.31 51.15